Math test has a mean of 41 and standard deviation of 5.0. Find the corresponding z scores for:

I. a test score of 18
II. a test score of 55

Please show all of your work.

Answer by stanbon(75887)

Math test has a mean of 41 and standard deviation of 5.0. Find the corresponding z scores for:
I. a test score of 18
z(18) = (18-41)/5 = -23/5 = -4.6
========================================

II. a test score of 55
z(55) = (55-41)/5 = 14/5 = 2.8
==================================
Cheers,
Stan H.